OpenStudy (anonymous):

Binomial expansion.. (X squared -2y) to the 5th

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[(x^2-2y)^5\] confirm please

OpenStudy (anonymous):

\[(x^2-y)^5= x^{10}-10x^8y+40x^6y^2-80x^4y^3+80x^2y^4-32y^5\]

OpenStudy (anonymous):

that's your binomial expansion

OpenStudy (anonymous):

ok how did you start off with X to the 10th

OpenStudy (anonymous):

ok have u heard of pascal's triangle?

OpenStudy (anonymous):

yes i have

OpenStudy (anonymous):

and the 5th would be 1 5 10 10 5 1 right?

OpenStudy (anonymous):

\[(x^2)^5\]

OpenStudy (anonymous):

yes u are write but look at the above...

OpenStudy (anonymous):

oh you multiply

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

u have to put the whole term to the highest power...

OpenStudy (anonymous):

ok but one like for the others you multiply them as well cause i remeber they said something about raising them

OpenStudy (anonymous):

yes u raise one power and reduce the other for the other term

OpenStudy (anonymous):

i was talking about the first term

OpenStudy (anonymous):

ok but would that be the final answer cause i saw my teacher do something else im not sur though

OpenStudy (anonymous):

well do u want the detailed solution?

OpenStudy (anonymous):

plugging in the binomial general formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i will show u how to do that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(a+b)^n=\sum_{k=0}^{n}\left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k}b^k\]

OpenStudy (anonymous):

wow well that looks a bit confusing

OpenStudy (anonymous):

hehehehe

OpenStudy (anonymous):

ok i will use the normal way

OpenStudy (anonymous):

ok please

OpenStudy (anonymous):

\[(a+b)^n=a^n+C^{n}_1a^{(n-1)}b+C^{n}_2a^{(n-2)}b^2+...+C^{n}_{n-1}ab^{(n-1)+b^n}b\]

OpenStudy (anonymous):

do u know this form...

OpenStudy (anonymous):

in this picture instead of a you plug in x^2, and instead of b you plug in -y

OpenStudy (anonymous):

ok well im gonna try working on it, thanks.